OpenStudy (anonymous):

how do i find out what (1/8)^(4/3) is equal to without using a calculator, i know it's 1/16, but how do i do it?

OpenStudy (anonymous):

1/cuberoot(8)^4

OpenStudy (anonymous):

the cube root of 8 is 2 2 to the fourth power is 16

OpenStudy (anonymous):

thank you very much!!

OpenStudy (callisto):

1/8 = (1/2)^3 (1/8)^(4/3) = [(1/2)^3] ^(4/3) = (1/2) ^(3x4/3) = (1/2)^4 = 1/16

OpenStudy (phi):

know your powers of 2

OpenStudy (callisto):

8 = 2x2x2 = 2^3